OpenStudy (anonymous):

Which expression is equivalent to (7^3)-^2

OpenStudy (anonymous):

I am guessing that you mean: (7^3)^(-2)? Since x^a^b = x^(a*b), you get 7^-6 Since x^(-a) = 1/(x^a), you get 1/(7^6) 7^6 is equal to 117649, so your answer would be: 1/117,649

OpenStudy (anonymous):

yep here are the chosies 1 over 7*7*7*7*7*7 7 1 over 6*6*6*6*6*6*6* 1 over 7

OpenStudy (anonymous):

The answer would be 1 over 7*7*7*7*7*7 since that is the same as 1/7^6 =)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

fan me and i will back